How many minutes would be saving, by making a 100 kilometer trip at 125 km/hr instead of 80 hr/hr?
Answer by stanbon(57361)

How many minutes would be saving, by making a 100 kilometer trip at 125 km/hr instead of 80 km/hr?
--------------
Faster trip time:
distance = 100 km ; rate = 125 km/hr ; time = 100/125 hr = 4/5 hr
-------------------
Slower trip time:
distance = 100 km ; rate = 80 km/hr ; time = 100/80 hr = 5/4 hr
----------------------------
Time difference:
5/4 - 4/5 = (25-16)/20 = 9/20 hr
-------------
Convert to minutes:
(9/20) hr = (9/20)(60 min) = 27 minutes
========================================
Cheers,
Stan H.